# tableTAB. 0

### v2

@@ II Plasmons in "strange" metals
"Strange" metals are at present one of the most interesting research fields in solid state physics [1]. Due to the strong interaction between their charge carriers, they show a deviation from the Fermi-liquid behaviour, e.g., they do not show a quadratic but a linear temperature dependence of the electrical resistivity or linear in energy scattering rate in Angle-Resolved Photoemission-Spectroscopy (ARPES)] [2, 3]. Moreover, there is no saturation at the Mott-Ioffe-Regel limit [4]. The unconventional and in some cases high-temperature superconductivity detected in these materials is supposed to be related to their "strange" normal-state electronic structure. Doped cuprates are prototypes of these "strange" metals. The non-Fermi liquid properties could be explained by a continuum of excitations up to an ultraviolet cutoff frequency $\omega_{\text{c}}$ in the electronic susceptibility [5], leading to a phenomenological marginal Fermi liquid theory. Integrating over this continuum yields a linear in energy imaginary part of the self-energy or scattering rate [6]. Electron Energy-Loss Spectroscopy (EELS) is a suitable experimental method to verify the existence of such a continuum because it measures the imaginary part of the electronic susceptibility $\Im[\chi({\bf q},\omega)]$. Here ${\bf q}$ is the momentum and $\omega$ is the energy. In nearly-free electron metals, there exists a $2k_{\text{F}}$ wide stripe of a continuum of particle-hole excitations [7] which starts at $q_{\mathrm{crit}}\approx\omega/v_{\text{F}}$, where $v_{\text{F}}$ is the Fermi velocity. In simple metals collective excitations (plasmons) exist at the order of the Fermi energy. At higher energies they are (Landau) damped by a decay into particle-hole excitations above the critical momentum: $q_{\mathrm{crit}}\approx\omega_{\text{P}}/v_{\text{F}}$. Usually, $q_{\mathrm{crit}}$ is between half and one Å-1.
Surprisingly, early transmission EELS (T-EELS) studies of the highly correlated doped cuprates, using dedicated T-EELS spectrometers [8, 9, 10] with high momentum resolution, showed this behavior with a weak damping of plasmons below $q_{\mathrm{crit}}$ [11, 12, 13, 14, 15, 16]. On the other hand, T-EELS studies using transmission electron microscopes with weak momentum resolution [17, 18] showed no plasmon at 0.3 Å-1 but a continuum. The difference can be easily explained by the poor momentum resolution in the TEM experiments ($\Delta q=$ 10 and 30 Å-1) which averages over the whole Brillouin zone (BZ) and therefore measures predominantly the Landau continuum above $q_{\mathrm{crit}}$.
@@
In this context, we mention that in various cuprates, well pronounced dispersive acoustic plasmons were detected by RIXS [29, 30, 31]. In all these measurements, weakly damped plasmons were detected for momentum ranging up to half of the size of the BZ.
@@ III Plasmon dispersion in layered systems
The calculation of the susceptibility of the many-body system of the charge carriers in solids is a challenging task. The susceptibility $\chi_{0}$ for a non-interacting one-band electron liquid is given by [43]:
$\chi_{0}(\omega,{\bf q})=\int_{\mathrm{BZ}}A({\bf q,k})\frac{2F({\bf k})\Delta E({\bf q,k})}{(\omega+i\Gamma)^{2}-\Delta E^{2}({\bf q,k})}d^{3}k\,.$ (2)
Here $\Delta E=E_{\bf k+q}-E_{\bf k}$, $E_{\bf k}$ are the band energies of the electrons having a momentum $\bf k$, $A({\bf q,k})$ is related to matrix elements, $\Gamma$ is the lifetime broadening of the particle-hole excitations, and $F({\bf k})$ is the Fermi function.
While $\chi_{0}$ is the Lindhard susceptibility for single-particle excitations related to an external field, stemming from the field of the scattering electron, $\chi$ is the susceptibility for the total field, including the induced one. Running a self-consistency cycle, we obtain the mean field RPA result:
@@
Here, V(q) is the Fourier transformed Coulomb interaction between the charge carriers and $\epsilon_{\text{b}}$ is the background dielectric constant.
@@
$\omega_{\mathrm{P}}(0)^{2}=\frac{4\pi Ne^{2}}{\epsilon_{\text{b}}m^{*}}$ (4)
with $N$ being the density of the charge carriers, $m^{*}$ the effective mass, and $e$ the elementary charge.
@@
$\omega_{\mathrm{P}}(q)=\omega_{\mathrm{P}}(0)+A_{\mathrm{RPA}}q^{2}+.....;\hskip 14.22636ptA_{\mathrm{RPA}}=\frac{\hbar}{m}(\alpha_{1}+\alpha_{2}).$ (5)
$\alpha_{1}$ is determined by the finite compressibility of the electron liquid and is always positive. $\alpha_{2}$, which is always negative, is proportional to the size of the effective mass [26]. For free-electron metals only
$\alpha_{1}=\frac{m}{\hbar\omega_{\text{P}}(0)}\frac{3}{10}\left<v_{\mathrm{F}}^{2}\right>_{\bf q}$ (6)
@@
For metals where the band dispersion is strongly reduced by a finite effective mass $m^{*}$, the negative $\alpha_{2}$ may dominate the optical plasmon dispersion, leading in total to a negative dispersion [26].
When the plasmon merges into the single-particle continuum, it decays into particle-hole excitations. This process is termed Landau damping.
@@
where $q_{||}~{}(q_{\perp})$ is the momentum parallel (perpendicular) to the layer [38]. Therefore, in a layered system, the plasmon dispersion is not only determined by the compressibility of the electron liquid (see Eq. 5) but also by the structure factor $V({\bf q)}$:
$\omega_{\mathrm{P}}(q)^{2}=s^{2}q_{||}^{2}+\omega_{\mathrm{P}}(0)^{2}\frac{q_{||}d}{2}\frac{\mathrm{sinh}(q_{||}d)}{\mathrm{cosh}(q_{||}d)-\mathrm{cos}(q_{\perp}d)}$ (10)
@@ IV EELS data
T-EELS measurements were performed on single-crystalline Sr2RuO4 layers at room temperature. Fig. 1 shows the crystal structure of Sr2RuO4 which consists of RuO2 layers stacked with SrO spacer layers along the ${\bf c}$-axis direction.
@@
The thin films were characterized by in-situ electron diffraction and the crystallographic axes were oriented with respect to the $({\bf a,c})$ plane [see Fig. 2 (a)]. The distance $d=$ 6.36 Å  between the RuO2 layers is half the $\bf c$-axis lattice constant. The thin lamellas for T-EELS with a normal direction parallel to the ${\bf b}$-axis were cut with a focused ion beam. In this way, EELS experiments were possible with the momentum parallel to the ${(\bf a,c})$ plane. We emphasize that this momentum range is different from our previous EELS study on Sr2RuO4 [36] where we covered the momentum range in the ${(\bf a,b})$ plane.
T-EELS was performed with a primary electron energy of 80 keV and with an energy and momentum resolution of 120 meV and 0.04 Å-1, respectively. The momentum-resolved EELS data were sequentially recorded in the ($q_{||}=q_{\text{a}},q_{\perp}$) momentum plane as depicted in Fig. 2(b). In Fig. 3 we show typical EELS intensities as a function of the energy for various $q_{||}$ and $q_{\perp}$ values. With increasing $q_{||}$ the plasmon energy slightly increases, whereas for increasing $q_{\perp}$ the plasmon energies decrease. At low total momentum it is difficult to see a well defined plasmon due to the high intensity of the quasi-elastic peak. The same holds for high total momentum because the T-EELS cross section is decreasing with $1/q^{2}$ (see below). Except for the described cases, well developed and dispersing plasmon excitations below 1.8 eV are visible. From the fit to the data we obtain the energy, the width, and the intensity of the plasmon excitations (see Methods).
@@
The difference between RPA dispersion and experimental data for $q_{||}\geq q_{\mathrm{crit}}$ can be explained by Landau damping. For these wave vectors, the plasmon decays into the particle-hole excitations and the spectral weight is caused by the Lindhard continuum. This leads to a reduction of the energy of the maximum in the loss function. This reduction has been observed in our calculations presented in Ref. [36]. A similar behavior has been also observed for Al [48]. In Fig. 4 we also show the continuum of the single-particle intra-band transitions $\chi_{0}$ (see the grey region) calculated using an unrenormalized TB band structure [47] (see Methods). The optical plasmon merges into the continuum near $q_{||}\approx$ 0.4 Å-1. We assign this value to the critical momentum $q_{\mathrm{crit}}$. This value is further corroborated by the rapid increase of the plasmon width at this wave vector (Fig. 5).
@@ V Optical and acoustic plasmons in layered "strange" metals
Well defined plasmon exist in the complete range which is not covered by the range of single-particle particle excitations calculated in the mean-field RPA theory. Optical plasmons exist in $\approx$ 15 % of the BZ . The rest is determined by a continuum of intra-band single particle excitations, which strongly dampen the plasmon excitation in excellent agreement with previous studies [36] (see Fig. 5). There is no sign of a reduction of the coherent plasmon range due to an over-damping discussed in terms of holographic theories [19, 20].
@@
Figure 4: Plasmon dispersion along the momentum $q_{||}$ parallel to the layers for several $q_{\perp}$ perpendicular to the layers (squares) together with calculations within the framework of the Fetter model (solid lines). The $q_{\perp}$ values 0, 0.1, 0.2, 0.3, and 0.4 Å-1 are marked by purple, blue, green, yellow, and red colour symbols, respectively (see Fig. 3). We have added also the data from the optical plasmon dispersion derived in our previous publication (dark purple stars) [36]. The horizontal error bars originate from the finite momentum resolution while the vertical ones are related to the finite spectral resolution of the EEL spectrometer (the fitting error is small in comparison). The region marked in grey corresponds to the susceptibility $\chi_{0}$ calculated from a tight binding band structure (see Methods). The excitations in the continuum range are marked by open symbols. The colored dashed curves represent theoretical curves considering finite momentum resolution, i.e., integration within the EEL collection aperture.
In the following we compare the present EELS results with theoretical predictions in a more quantitative way. In our previous publication we analyzed the plasmon dispersion by a calculation of the susceptibility together with an adjustment of the strength of the Coulomb potential to $\omega_{\mathrm{P}}(0)$. The comparison with the experimental dispersion yields good agreement when using an unrenormalized TB band structure ($\frac{m^{*}}{m_{0}}=1$) instead of a renormalized band structure with an effective mass $\frac{m^{*}}{m_{0}}\approx 4.5$ taken from angle-resolved photoemission spectroscopy (ARPES) experiments [49]. In this quite indirect analysis, we concluded that the plasmon dispersion is not influenced by the mass renormalization observed by ARPES.
Here we perform a similar, but much more direct analysis. Fitting the dispersion of the optical plasmon dispersion for $q\leq q_{\mathrm{crit}}\approx$ 0.4 Å-1 (see Fig. 4 purple data) with Eq. (1) we derive a dispersion coefficient $A=$ 1.8 $\pm$ 0.8 (eVÅ)2 which is in agreement with the more precise result from our previous publication $A=$ 2.1 $\pm$ 0.2 (eVÅ)2. Using the unrenormalized tight binding band structure [47] we receive a squared Fermi velocity $v_{\text{F}}^{2}\right>_{100}=$ 3.4 (eVÅ)2 and the dispersion coefficient $A=$ 2.5 (eVÅ)2. The latter value is close to the experimental value. On the other hand, using the effective mass $\frac{m^{*}}{m_{0}}\approx$ 4.5 from ARPES [49] we derive a negative dispersion with $A=$0.012 (eVÅ)2 which is at variance with the experimental data. This effective mass dependence of the plasmon dispersion was already discussed in our previous paper [36]. Thus our present and the previous experimental data of the dispersion of the optical plasmon indicates that the long wavelength dispersion can be explained in the framework of a mean-field RPA theory.
In Fig. 5 we present the plasmon width as a function of $q_{||}$. We show data derived from Fig. 3 and from our previous EELS experiments [36]. Within error bars there is a good agreement between the two datasets. The plasmon width at zero momentum is smaller than the plasmon energy, indicating a coherent collective charge excitation. The width below $q_{||}=0.3$-1 is constant. Near $q_{||}=$0.4 Å-1 the width increases, indicating the merging of the plasmon dispersion into the single particle continuum at a $q_{\mathrm{crit}}\approx$ 0.4 Å-1. (See also Fig. 4). As the calculations for the continuum are performed without band-renormalisation this indicates again that at the relative high plasmon energy, the effective mass is close to one and supports the formation of resilient quasi-particles [46].
@@
Figure 5: Plasmon width at half maximum in dependence on the momentum transfer parallel to the layers at $q_{\perp}=0$. For comparison the corresponding data from a Drude fit of our previous EELS measurements reported in Ref.  [36] was added as dark purple stars. The excitations in the continuum range are marked by open symbols.
@@
In Fig. 6 we depict the plasmon intensities as a function of $q_{||}$ and $q_{\perp}$ (see Methods section for details of the evaluation). The decay of the total spectral weight of the plasmon resonances at large momentum transfers is approximately proportional to $q^{-2}$ as observed for conventional bulk plasmons and predicted by the longitudinal f-sum rule and the theoretical dielectric function of the Fetter model (see Methods). The decay observed for $q_{\perp}\neq 0$ in the long wavelength limit is also consistent with the theoretical dielectric response as well as a version of the long-wavelength sum rule implying that the full spectral weight at $q=0$ is concentrated in the bulk plasmon $q_{\perp}=0$ mode (see Methods and Ref. [50] for the sum rules). The shift of the maximal integrated loss intensity towards smaller momentum transfer predicted by theory may be attributed to shortcomings of the Fetter model dielectric function (see Methods), impact of the zero loss, and limited experimental resolution.
@@
In the following we discuss the acoustic plasmon data (see red data and line in Fig. 4). Despite the gap at low energy due to a finite energy resolution, the dispersion extrapolates to zero energy typical of an acoustic plasmon. The derived experimental plasmon velocity is $v_{\mathrm{P}}\approx$ 4.7 eVÅ. From Eq. (11) we derive $v_{\mathrm{P}}\approx$ 4.8 eVÅ   in very good agreement with the experimental value. This indicates that the first term in Eq. (11) due to the finite Fermi velocity is small compared to the term which only depends on $\omega_{\mathrm{P}}(0)$ and $d$. Thus, for a given $\omega_{\mathrm{P}}(0)$ the acoustic plasmon dispersion only depends on $d$ and is not influenced by a large Fermi velocity but hints to a reduced one due to correlation effects (enhanced effective mass). Unfortunately, the finite energy resolution in the present T-EELS experiment does not allow a quantitative determination of the effective mass at low energies. However, RIXS data on cuprates indicate, that an enhanced effective mass is necessary to describe the plasmon velocity of the acoustic plasmons.

### v3

@@ II Plasmons in "strange" metals
"Strange" metals are at present one of the most interesting research fields in solid state physics [1]. Due to the strong interaction between their charge carriers, they show a deviation from the Fermi-liquid behaviour, e.g., they do not show a quadratic but a linear temperature dependence of the electrical resistivity or linear in energy scattering rate in Angle-Resolved Photoemission-Spectroscopy (ARPES)] [2, 3]. Moreover, there is no saturation at the Mott-Ioffe-Regel limit [4]. The unconventional and in some cases high-temperature superconductivity detected in these materials is supposed to be related to their "strange" normal-state electronic structure. Doped cuprates are prototypes of these "strange" metals. The non-Fermi liquid properties could be explained by a continuum of excitations up to an ultraviolet cutoff frequency $\omega_{\text{c}}$ in the electronic susceptibility [5], leading to a phenomenological marginal Fermi liquid theory. Integrating over this continuum yields a linear in energy imaginary part of the self-energy or scattering rate [6]. Electron Energy-Loss Spectroscopy (EELS) is a suitable experimental method to verify the existence of such a continuum because it measures the imaginary part of the electronic susceptibility $\Im[\chi({\bf q},\omega)]$. Here ${\bf q}$ is the momentum and $\omega$ is the energy. In nearly-free electron metals, there exists a $2k_{\text{F}}$ (with $k_{\text{F}}$ equal to the Fermi wave vector) wide stripe of a continuum of particle-hole excitations [7] which starts at $q\approx\omega/v_{\text{F}}$, where $v_{\text{F}}$ is the Fermi velocity. In simple metals collective excitations (plasmons) exist below the critical wave vector $q_{\mathrm{crit}}$ $\approx\omega_{\text{P}}/v_{\text{F}}$ which is determined by the plasmon energy $\omega_{\text{P}}$ and the Fermi velocity $v_{\text{F}}$. Above this momentum the plasmons merge into the continuum and therefore are (Landau) damped by a decay into particle-hole excitations. Usually, $q_{\mathrm{crit}}$ is between half and one Å-1.
Surprisingly, early transmission EELS (T-EELS) studies of the highly correlated doped cuprates, using dedicated T-EELS spectrometers [8, 9, 10] with high momentum resolution, showed this behavior with a weak damping of plasmons below $q_{\mathrm{crit}}$ [11, 12, 13, 14, 15, 16]. On the other hand, T-EELS studies using transmission electron microscopes with weak momentum resolution [17, 18] showed no plasmon but a continuum. The difference can be easily explained by the poor momentum resolution in the TEM experiments ($\Delta q=$ 10 and 30 Å-1) which averages over the whole Brillouin zone (BZ) and therefore measures predominantly the Landau continuum above $q_{\mathrm{crit}}$.
@@
In this context, we mention that in various cuprates, well pronounced dispersive acoustic plasmons were detected by resonant inelastic X-ray scattering (RIXS)  [29, 30, 31]. In all these measurements, weakly damped plasmons were detected for momentum ranging up to half of the size of the BZ.
@@ III Plasmon dispersion in layered systems
The calculation of the Lindhard-Ehrenreich-Cohen susceptibility of the many-body system of the charge carriers in solids is a challenging task. The susceptibility $\chi_{0}$ for a non-interacting one-band electron liquid is given by [43, 14]:
$\chi_{0}(\omega,{\bf q})=\int_{\mathrm{BZ}}M({\bf q,k})\frac{2F({\bf k})\Delta E({\bf q,k})}{(\omega+i\Gamma)^{2}-\Delta E^{2}({\bf q,k})}d^{3}k\,.$ (2)
Here $\Delta E=E_{\bf k+q}-E_{\bf k}$, $E_{\bf k}$ are the band energies of the electrons having a momentum $\bf k$, $M({\bf q,k})$ is related to matrix elements, $\Gamma$ is the lifetime broadening of the particle-hole excitations, and $F({\bf k})$ is the Fermi function.
While $\chi_{0}$ is the Lindhard-Ehrenreich-Cohen susceptibility for single-particle excitations related to an external field, stemming from the field of the scattering electron, $\chi$ is the susceptibility for the total field, including the induced one. Running a self-consistency cycle, we obtain in the mean field RPA the result:
@@
Here, V(q) is the Fourier transformed Coulomb interaction between the charge carriers and $\epsilon_{\text{b}}$ is the background dielectric constant. The dielectric function can be calculated from
$\epsilon({\bf q},\omega)=\epsilon_{b}-V({\bf q})\chi_{0}({\bf q},\omega).$ (4)
$\epsilon_{\text{b}}$ is the background dielectric function.
@@
$\omega_{\mathrm{P}}(0)^{2}=\frac{4\pi Ne^{2}}{\epsilon_{\text{b}}(m^{*}/m_{0})}$ (5)
with $N$ being the density of the charge carriers, $m^{*}$ the effective mass, $m_{0}$ the unrenormalised mass, and $e$ the elementary charge.
@@
$\omega_{\mathrm{P}}(q)=\omega_{\mathrm{P}}(0)+A_{\mathrm{RPA}}q^{2}+.....;\hskip 14.22636ptA_{\mathrm{RPA}}=(A_{1}+A_{2}).$ (6)
$A_{1}$ is related by the finite compressibility or the squared averaged Fermi velocity of the electron liquid and is always positive. $A_{2}$, which is always negative, is proportional to the size of the effective mass [26].
For free-electron metals only
$A_{1}=\frac{1}{\omega_{\text{P}}(0)}\frac{3}{10}\left<v_{\mathrm{F}}^{2}\right>_{\bf q}$ (7)
@@
For metals where the band dispersion is strongly reduced by a finite effective mass $m^{*}/m_{0}$, the negative $A_{2}$ may dominate the optical plasmon dispersion, leading in total to a negative dispersion [26].
@@
where $q_{||}~{}(q_{\perp})$ is the momentum parallel (perpendicular) to the layer [38]. Therefore, in a layered system, the plasmon dispersion is not only determined by the compressibility and by the effective mass of the electron liquid (see Eq. 6) but also by the structure factor $V({\bf q)}$, depending on $q_{||}d$ and $q_{\bot}d$
$\omega_{\mathrm{P}}(q)^{2}=s^{2}q^{2}+\omega_{\mathrm{P}}(0)^{2}\frac{q_{||}d}{2}\frac{\mathrm{sinh}(q_{||}d)}{\mathrm{cosh}(q_{||}d)-\mathrm{cos}(q_{\perp}d)}$ (11)
@@ IV EELS data
T-EELS measurements were performed on single-crystalline Sr2RuO4 at room temperature. Fig. 1 shows the crystal structure of Sr2RuO4 which consists of RuO2 layers stacked with SrO spacer layers along the ${\bf c}$-axis direction.
@@
The thin films were characterised by in-situ electron diffraction and the crystallographic axes were oriented with respect to the $({\bf a,c})$ plane [see Fig. 2 (a)]. The distance $d=$ 6.36 Å  between the RuO2 layers is half the $\bf c$-axis lattice constant. The thin lamellas for T-EELS with a normal direction parallel to the ${\bf b}$-axis were cut with a focused ion beam. In this way, EELS experiments were possible with the momentum parallel to the ${(\bf a,c})$-plane. We emphasize that this momentum range is different from our previous EELS study on Sr2RuO4 [36] where we covered the momentum range in the (${\bf a}$,${\bf b}$)-plane.
T-EELS was performed with a primary electron energy of 80 keV and with an energy and momentum resolution of 120 meV and 0.04 Å-1, respectively. The momentum-resolved EELS data were sequentially recorded in the ($q_{||}=q_{\text{a}},q_{\perp}$) momentum plane as depicted in Fig. 2(b). In Fig. 3 we show typical EELS intensities as a function of the energy for various $q_{||}$ and $q_{\perp}$ values. With increasing $q_{||}$ the plasmon energy slightly increases, whereas for increasing $q_{\perp}$ the plasmon energies decrease. At low total momentum it is difficult to see a well defined plasmon due to the high intensity of the quasi-elastic peak. The same holds for high total momentum because the T-EELS cross section is decreasing with $1/q^{2}$ (see below). Except for the described cases, well developed and dispersing plasmon excitations below 1.8 eV are visible. From the fit to the loss data with a Drude function, we obtain the energy, the width, and the intensities (see Methods).
@@
The difference between extrapolated RPA plasmon dispersion and experimental data for $q_{||}\geq q_{\mathrm{crit}}$ (see Fig. 4) can be explained by Landau damping. For these wave vectors, the plasmon decays into over-damped plasmon excitations and into spectral weight which is caused by the Lindhard continuum. This leads to a reduction of the energy of the maximum in the total loss function. This reduction has been observed in our calculations presented in Ref. [36]. A similar behavior has been also observed for Al [48]. In Fig. 4 we also show the continuum of the single-particle intra-band transitions $\chi_{0}$ (see the grey region) calculated using an unrenormalised TB band structure [47] (see Methods). The optical plasmon merges into the continuum near $q_{||}\approx$ 0.4 Å-1. We assign this value to the critical momentum $q_{\mathrm{crit}}$. This value is further corroborated by the rapid increase of the plasmon width at this wave vector (see Fig. 6).
@@ V Optical and acoustic plasmons in layered "strange" metals
Well defined plasmons exist in the complete range which is not covered by the range of single-particle particle excitations calculated in the mean-field RPA theory. Optical plasmons exist in $\approx$ 15 % of the BZ . The rest is determined by a continuum of intra-band single particle excitations, which strongly dampen the plasmon excitation in excellent agreement with previous studies [36]. There is no sign of a reduction of the coherent plasmon range due to an over-damping discussed in terms of holographic theories [19, 20].
Fitting the dispersion of the optical plasmon dispersion for $q_{||}\leq q_{\mathrm{crit}}\approx$ 0.4 Å-1 (see Fig. 4 purple data) with Eq. (1) we derive a dispersion coefficient $A=$ 1.8 $\pm$ 0.8 eVÅ2 which is in agreement with the more precise result from our previous publication [36] $A=$ 2.1 $\pm$ 0.2 eVÅ2. In the following we discuss the latter $A$ value from our previous publication.
@@
Figure 4: Plasmon dispersion along the momentum $q_{||}$ parallel to the layers for several $q_{\perp}$ perpendicular to the layers (squares) together with calculations within the framework of the Fetter model (solid lines). The $q_{\perp}$ values 0, 0.1, 0.2, 0.3, and 0.4 Å-1 are marked by purple, blue, green, yellow, and red colour symbols, respectively (see Fig. 3). We have added also the data from the optical plasmon dispersion derived in our previous publication dark purple stars) [36]. The horizontal error bars originate from the finite momentum resolution while the vertical ones are related to the finite spectral resolution of the EEL spectrometer ( the fitting error is small in comparison). The region marked in grey corresponds to the susceptibility $\chi_{0}$ calculated from a tight binding band structure (see below and Methods). The excitations in the continuum range are marked by open symbols. The colored dashed curves represent theoretical curves considering finite momentum resolution, i.e., integration within the EEL collection aperture.
Figure 5: Calculated imaginary part of the susceptibility $\chi_{0}(\omega,q_{||},q_{\perp}=0)$ (second column), loss function $\Im\left\{-\frac{1}{\epsilon(\omega,q_{||},q_{\perp}=0)}\right\}$ (third column), and plasmon dispersion (fourth column) along the momentum $(q_{||},q_{\perp}=0)$ together with least squares fit (red lines) in the momentum range $q_{||}$= 0 to 0.4 Å-1. In the calculation various effective masses were used: first row $m^{*}/m_{0}=$1, second row $m^{*}/m_{0}=$3.5, and an energy dependent effective mass $m^{\ast}(\omega)/m_{0}$ shown in the third row. The red labels in the susceptibility in the first row are related to the origin of the three electronic bands in Sr2RuO4.
@@
Figure 6: Plasmon width at half maximum in dependence on the momentum transfer parallel to the layers at $q_{\perp}=0$. For comparison the corresponding data from a Drude fit of our previous EELS measurements reported in Ref.  [36] was added as dark purple stars. The energy of the maximum of the excitations in the continuum range (above $q_{||}$ are marked by open symbols.
Next we compare the present EELS data of the optical plasmon dispersion with those derived from theoretical calculations of the loss function. In Fig. 5 we present these results for the susceptibility, the loss function, and the optical plasmon dispersion for an unrenormalised band structure [47] ($m^{*}/m_{0}=$1), a constant effective mass ($m^{*}/m_{0}=$ 3.5), and an energy dependent effective mass ranging from 3.5 at low energies to 1 above $\approx$ 0.2 eV. The latter was taken from optical spectroscopy [46] together with an extrapolation to 1 at higher energies.
Using the unrenormalised tight binding band structure [47] we receive below $q_{\text{crit}}$ the dispersion coefficient $A=$ 2.8 eV Å2. The latter value is close to the experimental value $A=2.1\pm 0.2$ eVÅ2. The small difference between the experimental and the theoretical $A$ value can be explained by a finite thickness of the RuO2 layers which reduces the nominal $d$ to $d_{\mathrm{reduced}}=$ 4. This reduction of $d$ brings the theoretical $A$ value very close to the experimental one. Moreover, our calculations show that an enlargement of the half width with $\Gamma$ from about 0.1 to the experiment reduces the dispersion coefficient $A$ by $\approx 0.2$ eVÅ2.
On the other hand, using the constant effective mass $m^{*}/m_{0}=$ 3.5 from low-energy/low temperature optical spectroscopy we derive a reduced dispersion with $A=$ 1.1 eVÅ2 which is at variance with the experimental data. This effective mass dependence of the plasmon dispersion was already discussed in our previous paper [36]. We also remark that the susceptibility is strongly reduced in energy compared to the plasmon energy. Therefore in case of $m^{*}/m_{0}=$ 3.5 the plasmon is not merging into a continuum and exists well above $q_{\mathrm{crit}}$.
Using the energy dependent effective mass from optical spectroscopy below $\omega=$ 0.2 eV plus an extrapolation to $m^{*}(\omega)/m_{0}=1$ at higher energies we see the expected strong renormalization and an intensity increase in the low energy/momentum range. However, at higher energy/momentum the susceptibility is similar to that calculated with an unrenormalised band structure. This causes a plasmon dispersion which is very close to an unrenormalised dispersion. Thus our present and the previous experimental data of the dispersion of the optical plasmon indicates that the long wavelength dispersion can be explained in the framework of a mean-field RPA theory using an effective mass of one. Because for the cuprates an energy dependent effective mass is expected as well, probably the present result can be also explain the unrenormalised plasmon dispersion detected in the cuprates [14, 26].
The calculations clearly demonstrate that the low energy renormalization of the susceptibility/optically conductivity does not transfer into the high-energy plasmon dispersion. On the other hand, the renormalised acoustic plasmon dispersion observed by RIXS at low energies in cuprates [49] can be explained in this framework.
In Fig. 6 we present the optical plasmon width as a function of $q_{||}$. We show data derived from Fig. 3 and from our previous EELS experiments [36]. Within error bars there is a good agreement between the two datasets. The plasmon width at zero momentum is smaller than the plasmon energy, indicating a coherent collective charge excitation. The width below $q_{||}\approx 0.4$-1 is constant. Near $q_{||}=$0.4 Å-1 the width increases, indicating the merging of the plasmon dispersion into the single particle continuum at a $q_{\mathrm{crit}}\approx$ 0.4 Å-1. (See also Fig. 4). As the calculations for the continuum are performed without band-renormalisation this indicates again that at the relative high plasmon energy, the effective mass is close to one and supports the formation of resilient quasi-particles [46]. Already in our previous paper [36], we have emphasised that the width is almost constant in the studied low momentum range. This means that it is not related to electron-electron interaction which would lead to a quadratic increase of the width as a function of momentum transfer [50]. The fact that the width is smaller than the energy signals that the excitation is not overdamped due to fluctuations in a quantum critical system, in contrast to theoretical predictions [19]. Rather, it is caused, as in most metallic systems, studied by T-EELS, by a decay into interband transitions [51, 52, 27, 53]. These interbands are induced by a back-folding of bands from the second to the first Brillouin zone by a finite pseudo-potential. In this way, in recent RIXS data on p and n-typed cuprates [49], the difference in accoustic plasmon width could be described in this framework.
@@
In Fig. 7 we depict the plasmon intensities as a function of $q_{||}$ and $q_{\perp}$ (see Methods section for details of the evaluation). The decay of the total spectral weight of the plasmon resonances at large momentum transfers is approximately proportional to $q^{-2}$ as observed for conventional bulk plasmons and predicted by the longitudinal f-sum rule and the theoretical dielectric function of the Fetter model (see Methods). The decay observed for $q_{\perp}\neq 0$ in the long wavelength limit is also consistent with the theoretical dielectric response as well as a version of the long wavelength sum rule implying that the full spectral weight at $q=0$ is concentrated in the bulk plasmon $q_{\perp}=0$ mode (see Methods and Ref. [54] for the sum rules). The shift of the maximal integrated loss intensity towards smaller momentum transfer predicted by theory may be attributed to shortcomings of the Fetter model dielectric function (see Methods), impact of the zero loss, and limited experimental resolution.
@@
In the following we discuss the acoustic plasmon data (see red data and line in Fig. 4). Despite the gap at low energy due to a finite energy resolution, the dispersion extrapolates to zero energy typical of an acoustic plasmon. The derived experimental plasmon velocity is $v_{\mathrm{P}}\approx$ 4.7 eVÅ. From Eq. (12) we derive $v_{\mathrm{P}}\approx$ 4.8 eVÅ   in very good agreement with the experimental value. This indicates that the first term in Eq. (12) due to the finite Fermi velocity is small compared to the term which only depends on $\omega_{\mathrm{P}}(0)$ and $d$. Thus, for a given $\omega_{\mathrm{P}}(0)$ the acoustic plasmon dispersion only depends on $d$ and is not influenced by a large Fermi velocity but hints to a reduced one due to correlation effects (enhanced effective mass). Unfortunately, the finite energy resolution in the present T-EELS experiment does not allow a quantitative determination of the effective mass at low energies. However, low-energy RIXS data on cuprates indicate, that an enhanced effective mass is necessary to describe the plasmon velocity of the acoustic plasmons [49].